To decrease the pressure of the gas by 10% at constant temperature then change in volume should be????

Grade:12th pass

1 Answers

Arun
5 years ago
For an ideal gas at constant temperature, the volume is inversely proportional to pressure, so P1·V1 = P2·V2. With P2 = 0.90 P1, then V2 = V1/0.90, so ∆V = (1/0.90 – 1)V1 = 0.111 V1 = V1 / 9.